User Tools

Site Tools


handover:c1712_v1712

c1712/v1712

Hobart 26m

Module: HOB+0130/32000/1024, start pos 7462 GB

  • 0002UT - Experiment started ok. (Ellen)
  • 0007UT - previous source in this schedule not reached before new source was commanded - but not windstowed I don't think. (Ellen)
  • 0017UT - Windstowed, missed scans 343-0019 to 343-0041. (Ellen)
  • 0046UT - Windstowed, missed scans 343-0043 to 343-0211b. (Ellen)
  • 0143UT - pressure stablised at 4.2 microns Hg (Warren said to call at 5, Brett said 7), 70K stage at still varying between 50 and ~120K. Keeping an eye on them (Ellen)
  • 0221UT - Windstowed, missed scans 343-0221b to 343-0231a. (Ellen)
  • 0254UT - Windstowed, missed scans 343-0254 to 343-0346a. (Ellen)
  • 0255UT - 70K stage now constant at about 50K. (Ellen)
  • 2017.343.04:08:52.59;"Windstowed between 343-04:06:36 and 343-04:08:48
    2017.343.04:38:27.85;"Windstowed between 343-04:36:12 and 343-04:38:24
    2017.343.04:43:11.58;"Windstowed between 343-04:39:52 and 343-04:43:08
    2017.343.05:05:41.60;"Windstowed between 343-05:01:10 and 343-05:05:31
    2017.343.05:13:14.28;"Windstowed between 343-05:09:47 and 343-05:13:03
    2017.343.05:57:06.31;"Windstowed between 343-05:45:07 and 343-05:56:55
    2017.343.06:24:18.82;"Windstowed between 343-06:01:39 and 343-06:24:08
    2017.343.06:48:58.51;"Windstowed between 343-06:33:46 and 343-06:48:46
    2017.343.07:06:06.31;"Windstowed between 343-06:53:02 and 343-07:05:55
    2017.343.07:22:43.70;"Windstowed between 343-07:20:27 and 343-07:22:39
    2017.343.20:53:49.22;"Windstowed between 343.20:53:49 and 343.22:55:55
    2017.343.23:28:47.63;"Windstowed between 343-23:24:35 and 343-23:28:31
  • Antenna got half stuck in windstow between 2017.343.20:53:49.22–2017.343.22:55:55.59. Restarting schedule and windstow script fixed issue. No alarm.
  • New windstow script is messing up comments.py. To auto report windstows and missed scans for c1712ho, download the log to /vlbobs/ivs/logs/. This can be done using scp oper@hobart:/usr2/log/c1712ho.log /vlbobs/ivs/logs/ ,manually remove the line 2017.343.00:07:44.23;“Schedule resumed after end of wind stow.. Run python ask/comments/comments_hofix.py to generate the email. Run flogit as per normal but make sure you download a fresh copy of the log. You might have to delete another line if the telescope goes into stow before the exp end but doesnt recover before the exp finishes. Temp workaround until the old anemometer is fixed. (Arwin)

Katherine

Module: GSFC+012, start pos 3211 GB

  • 0002UT - Experiment started ok. (Ellen)
  • 0149UT - tsys in bands 09-12 consistently quite high (125-200). (Ellen)

Yarragadee

Module: HOB+01075, start pos 4167 GB

  • 0002UT - Experiment started ok. (Ellen)
  • 2017.343.06:10:00 UT windstow. The drives failed to recover, manually recovered at 2017.343.06:25:25, 'comments.py' will auto report this (Arwin).
  • Problem with VNC viewer for maser (14:35:30 UT), so far maserdelay and clkoff values are OK (mas)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/c1712_v1712.txt · Last modified: 2017/12/09 23:29 by Patrick Yates